abriskie:

Labor time = 4 1/2 hours Overhead rate = 95% Retail price of parts = $134.75 Total cost of job = $241.98 Hourly rate?

Allison:

Total cost is $241.98

Allison:

@Ultrilliam Am I right?

Ultrilliam:

I uh... Don't think so... .-.

Ultrilliam:

although I can't say for sure as I don't know how to calculate this stuff... where is psi! x.x

Allison:

241.98 - 134.75 = 107.23 107.25 / (4.5 x 1.95) = 107.25 / 8.78 = 12.22 hourly rate

Allison:

Your answer is 12.22 x.x

Ultrilliam:

okay that seems more likely x'D

Allison:

Labor : 12.22 x 4.5 = 54.99 Overhead: 54.99 x 95% = 52.24 Retail price of parts = 134.75 Total cost of job: = 241.98

abriskie:

thank you!

Allison:

Welcome :3

Allison:

Oh can you fan me? o:

Ultrilliam:

Welcome to QuestionCove

Allison:

f a i l

Ultrilliam:

don't mind the colors being off x'D

Allison:

Hm. Question Cove.

ravan364:

total cost is 241.98 fan me pls welcome to question cove

Allison:

Boi.

Ultrilliam:

Bruh

Allison:

The answer is 12.22.

ravan364:

no it is not alli

abriskie:

Allison is correct, thanks again.